I am working hard to get to the point where I could solve word problems in Algebra. Problem: The second angle of a triangular parking lot is four times as large as the first angle. The third angle is 45 degrees less than the sum of the other two angles. How long are the angles?
Answer by Abbey(339)

let the first angle = x
let the second angle =4x
let the third angle = x+4x-45
so when these are added together, they will total 180 degrees:
x+4x+x+4x-45=180
add 45 to both sides and combine like terms:
10x=225
x=22.5 degrees
so if the first angle =22.5 degrees, the second angle must = 90 degrees and the final angle equals 90+22.5-45 = 67.5 degrees.